How do you find sin(a+B), tan(a-B), cos(2B), and sin(a/2) if a and B are both in the first quadrant and sin(a)=2/3, and cos(B)=3/4?
Answer by Alan3354(69443)

How do you find sin(a+B), tan(a-B), cos(2B), and sin(a/2) if a and B are both in the first quadrant and sin(a)=2/3, and cos(B)=3/4?
-------------
Look at "Half angle formulas" in Wikipedia.